Freshman Seminar 21n: Elliptic Curve Projects

William Stein

April 15, 2003

Today you will give your second presentations. Your assignment is to send me an email (by Friday), with one paragraph about the other presentations you see today.

A very rough draft of your project is due next Tuesday, April 22. What this is will depend on the format you choose for your project. It could be a web page or printout.

Schedule

April 22 Third presentation (a rough draft is due)
April 29 Fourth presentation (a physical draft due)
May 6 Fifth presentation (get comments on your draft back)
May 13 Give a polished 40 minute presentation and final version

(Note: I will be gone May 5–6 (at IDA), but Grigor will lead class those two days.)